OpenStudy (anonymous):

Select one of the factors of 2x^2 + 16x – 18 x + 9 x + 1 x – 2 x – 8

OpenStudy (anonymous):

U know Factor Theorem?

OpenStudy (anonymous):

No I don't.

OpenStudy (anonymous):

If a polynomial f(x) is divided by (x-a) then (x-a) is a factor of f(x) if f(a)=0

OpenStudy (anonymous):

Does it help?

OpenStudy (mayankdevnani):

so as per factor therom you get x+1.

OpenStudy (mayankdevnani):

right @sauravshakya

OpenStudy (anonymous):

Nope... not x+1

OpenStudy (anonymous):

It is x+9

OpenStudy (anonymous):

I don't see how the factor theorem can help with this question (x^2 + 7x – 44) though?

OpenStudy (anonymous):

Did u get for 2x^2 + 16x – 18

OpenStudy (anonymous):

I got it but i'm not sure exactly how you worked it out i tried and got stuck

OpenStudy (anonymous):

f(x)=2x^2+16x-18 Now, U need to try for every options for which u will get f(a)=0 when it is divided by (x-a)

OpenStudy (anonymous):

In this case, When f(x) is divided by (x+9), f(-9)=2*(-9)^2+16*(-9)-18=0

OpenStudy (anonymous):

So, one of the factor of f(x) is (x+9)

OpenStudy (anonymous):

OK that makes sense.

OpenStudy (anonymous):

Now, can u do these type of problems?

OpenStudy (anonymous):

I think so thank you.

OpenStudy (anonymous):

welcome